Let C be the graph of y=1-x,-1sxsl. a) Write a vector-valued function r(t) that traces C. y b) Evaluate ca 11+4x2
Answer #1

y=1-2, IER51 Let a x=t; Vector valued function ges=1-t² reta (t, 1-22); rets1 riety = (1-2) -S I - I rectill at 1+ 41² +4E2 S